# ORCID: $0$-0002-6512-4229 (Om Prakash)

### v1

@@ 5 Constructions of quantum codes and comparison with the existing codes
The theory of quantum error-correcting codes plays a key role in quantum information theory. For a long time, it has been difficult to provide a satisfactory solution to the problem of protecting information from quantum noises. However, after the introduction of the first quantum error-correcting codes by Shor et al. Shor , a stream of great developments has emerged in information theory. Let $H_{q}(\mathbb{C})$ be a $q$-dimensional Hilbert vector space. Then the set of $n$-fold tensor product $H_{q}^{n}(\mathbb{C})=\underbrace{H_{q}(\mathbb{C})\otimes H_{q}(\mathbb{C})\otimes\cdots\otimes H_{q}(\mathbb{C})}_{n\rm\ times}$ is a $q^{n}$-dimensional Hilbert space. Here, a $q^{k}$ dimensional subspace of $H_{q}^{n}(\mathbb{C})$ is called a quantum code with parameters $[[n,k,d]]_{q}$ where $d$ is the minimum distance, and $k$ is the dimension of the quantum code. Also, $\mathcal{C}$ is dual-containing if $\mathcal{C}^{\perp}\subseteq\mathcal{C}$. Moreover, in $1997$, the quantum Singleton bound for binary codes was introduced by Knill and Laflamme Knill . In $1998$, Calderbank et al. Calderbank98 provided the quantum Singleton bound for all codes over finite fields as $k+2d\leq n+2.$ A quantum code is said to be a quantum MDS code if it attains the Singleton bound. In this section, we study quantum error correcting codes from $(\gamma,\Delta)$-cyclic code of length $n$ over $\mathscr{R}_{q,s}$. In order to construct quantum error-correcting codes, we first derive a necessary and sufficient condition for $(\gamma,\Delta)$-cyclic codes to be dual containing. Note that a quantum code $[[n,k,d]]_{q}$ is said to be better than $[[n^{\prime},k^{\prime},d^{\prime}]]_{q}$ if any one of the following or both hold:
@@
Now, we define a Gray map and study $\mathbb{F}_{q}$-images of $(\gamma,\Delta)$-cyclic codes. Let $GL_{s+1}(\mathbb{F}_{q})$ be the set of all $(s+1)\times(s+1)$ invertible matrices over $\mathbb{F}_{q}$. Now, $\varphi:\mathscr{R}_{q,s}\longrightarrow\mathbb{F}_{q}^{s+1}$ define by
@@
Proof
Let $a=(a_{0},a_{1},\dots,a_{n-1}),~{}b=(b_{0},b_{1},\dots,b_{n-1})\in\mathscr{R}_{q,s}^{n}$, where $a_{j}=\sum_{i=0}^{s}\zeta_{i}a_{j}^{i}$, $b_{j}=\sum_{i=0}^{s}\zeta_{i}b_{j}^{i}$ for $j=0,1,\dots,n-1$ and $a_{j}^{i},~{}b_{j}^{i}\in\mathbb{F}_{q}$. Then
@@
Moreover, $d_{G}(a,b)=\omega_{G}(a-b)=\omega_{H}(\varphi(a-b))=\omega_{H}(\varphi(a)-\varphi(b))=d_{H}(\varphi(a),\varphi(b))$. Hence, $\varphi$ is a distance preserving map.
Theorem 5.1
@@
Proof
Follows directly from Proposition 1 and the definition of the Gray map.
@@
Proof
Let $c=(c_{0},c_{1},\dots,c_{n-1})\in\mathcal{C}$ and $d=(d_{0},d_{1},\dots,d_{n-1})\in\mathcal{C}^{\perp}$ where $a_{j}=\sum_{i=0}^{s}\zeta_{i}c_{j}^{i}$, $b_{j}=\sum_{i=0}^{s}\zeta_{i}d_{j}^{i}$ for $j=0,1,\dots,n-1$ and $a_{j}^{i},~{}b_{j}^{i}\in\mathbb{F}_{q}$. Now, $c\cdot d=\sum_{j=0}^{n-1}c_{j}d_{j}=0$ gives $\sum_{j=0}^{n-1}(c^{0}_{j}d^{0}_{j}+c^{1}_{j}d^{1}_{j}+\dots+c^{s}_{j}d^{s}_{j})=0$. Again,
@@
Since $c\in\mathcal{C}$ and $d\in\mathcal{C}^{\perp}$ are arbitrary, $\varphi(\mathcal{C}^{\perp})\subseteq(\varphi(\mathcal{C}))^{\perp}$. On the other hand, as $\varphi$ is a bijective linear map, $\mid\varphi(\mathcal{C}^{\perp})\mid=\mid(\varphi(\mathcal{C}))^{\perp}\mid$. Therefore, $\varphi(\mathcal{C}^{\perp})=(\varphi(\mathcal{C}))^{\perp}$.
Lemma 2 (Grassl04 , Theorem 3)

### v2

@@ 5 Constructions of quantum codes and comparison with the existing codes
The quantum error-correcting codes play a pivotal role in quantum information theory. For a long time, it has been difficult to provide a satisfactory solution to the problem of protecting information from quantum noises. However, after the introduction of the first quantum error-correcting codes by Shor et al. [42], a stream of great developments has emerged in information theory. Let $H_{q}(\mathbb{C})$ be a $q$-dimensional Hilbert vector space. Then the set of $n$-fold tensor product $H_{q}^{n}(\mathbb{C})=\underbrace{H_{q}(\mathbb{C})\otimes H_{q}(\mathbb{C})\otimes\cdots\otimes H_{q}(\mathbb{C})}_{n\rm\ times}$ is a $q^{n}$-dimensional Hilbert space. Here, a $q^{k}$ dimensional subspace of $H_{q}^{n}(\mathbb{C})$ is called a quantum code with parameters $[[n,k,d]]_{q}$ where $d$ is the minimum distance, and $k$ is the dimension of the quantum code. Also, $\mathcal{C}$ is dual-containing if $\mathcal{C}^{\perp}\subseteq\mathcal{C}$. Moreover, in $1997$, the quantum Singleton bound for binary codes was introduced by Knill and Laflamme [28]. In $1998$, Calderbank et al. [11] provided the quantum Singleton bound for all codes over finite fields as $k+2d\leq n+2.$ A quantum code is said to be a quantum MDS code if it attains the Singleton bound.
In this section, we first briefly review the mathematical representation of the quantum states, the operators acting on these states, and then we construct quantum codes from $(\gamma,\Delta)$-cyclic codes over $\mathscr{R}_{q,s}$.
5.1 Quantum states and operators over qudits
For a quantum system with $\Gamma$ levels, the state of a unit system, a qudit, is a superposition of $\Gamma$ basis states of the system given by
|$\ket{\psi}_{\Gamma}$ $=\sum_{i=0}^{\Gamma-1}a_{i}\ket{i}_{\Gamma},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{ where }a_{i}\in\mathbb{C}\text{ and }\sum_{i=0}^{\Gamma-1}|a_{i}|^{2}=1,$
where the subscript $\Gamma$ refers to dimension of the unit quantum system. Also, |$\ket{\psi}_{\Gamma}=[a_{0}\leavevmode\nobreak\ a_{1}\leavevmode\nobreak\ \dots\leavevmode\nobreak\ a_{\Gamma-1}]^{\mathrm{T}}$ and |$\ket{i}_{\Gamma}=\mathbf{e}_{{(i+1)}}^{{(\Gamma)}}$, where $\mathbf{e}_{{(i+1)}}^{{(\Gamma)}}$ is a vector in $\mathbb{C}^{\Gamma}$ with the $(i+1)^{\mathrm{st}}$ element being $1$ and rest of the elements being 00.
From the second postulate of quantum mechanics, the operators acting on a quantum system belong to the unitary group $\mathrm{U}(\Gamma)$, which is a subset of $\mathbb{C}^{\Gamma\times\Gamma}$. As the cardinality of $\mathrm{U}(\Gamma)$ is infinite, we represent its elements in terms of a basis of $\mathbb{C}^{\Gamma\times\Gamma}$.
For $\Gamma=2$, the Pauli basis $\mathcal{P}$ is the popularly chosen unitary basis.
$\mathcal{P}\!=\!\left\{\!\mathrm{I}\!=\!\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\!,\mathrm{X}\!=\!\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\!,\mathrm{Y}\!=\!\begin{bmatrix}0&-\mathrm{i}\\ \mathrm{i}&0\end{bmatrix}\!,\mathrm{Z}\!=\!\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}\right\},$
where $\mathrm{i}=\sqrt{-1}$. The generalized version of the Pauli group for arbitrary $\Gamma$, known as the Weyl-Heisenberg group, is defined by
$\mathcal{G}_{\Gamma}^{(g)}=\left\{\omega_{\Gamma}^{l}\mathrm{X}_{\Gamma}(a)\mathrm{Z}_{\Gamma}(b)|a,b,l\in\mathbb{Z}_{\Gamma}\right\},$
where $\omega_{\Gamma}\!=\!\mathrm{e}^{\!\frac{\mathrm{i}2\pi}{\Gamma}}$, $\mathrm{X}_{\Gamma}(a)\!\ket{c}_{\Gamma}\!:=\!\!\ket{(a\!+\!c)\leavevmode\nobreak\ \mathrm{mod}\leavevmode\nobreak\ \Gamma}_{\Gamma}$, and $\mathrm{Z}_{\Gamma}(b)\!\ket{c}_{\Gamma}:=\omega_{\Gamma}^{bc}\!\ket{c}_{\Gamma}$ for every $c\in\mathbb{Z}_{\Gamma}$. The generalized Pauli basis $\mathcal{P}$ [34]
$\mathcal{G}_{\Gamma}=\left\{\mathrm{X}_{\Gamma}(a)\mathrm{Z}_{\Gamma}(b)|a,b\in\mathbb{Z}_{\Gamma}\right\}.$ (3)
is obtained by neglecting the phase $\omega_{\Gamma}^{l}$ in $\mathcal{G}_{\Gamma}$. The basis operator of the form $\mathrm{X}_{\Gamma}(a)\mathrm{Z}_{\Gamma}(b)$ is uniquely represented by a vector of length $2$ defined over ring $\mathbb{Z}_{\Gamma}$, namely $[a|b]_{\Gamma}$ as
$\mathrm{X}_{\Gamma}(a)\mathrm{Z}_{\Gamma}(b)\equiv[a|b]_{\Gamma}.$
Next, we define a trace operation over the field elements as follows:
Definition 5.1 ([27]).
The field trace $\mathrm{Tr}_{p^{m}/p}(\cdot)$ is an $\mathbb{F}_{p}$-linear function $\mathrm{Tr}_{p^{m}/p}:\mathbb{F}_{p^{m}}\rightarrow\mathbb{F}_{p}$, given by $\mathrm{Tr}_{p^{m}/p}(\kappa)=\sum_{i=0}^{m-1}\kappa^{p^{i}}$, where $\kappa\in\mathbb{F}_{p^{m}}$.
The function $\mathrm{Tr}_{p^{m}/p}(\cdot)$ is said to be $\mathbb{F}_{p}$-linear as $\mathrm{Tr}_{p^{m}/p}(a\kappa+b\chi)=a\hskip 1.9919pt\mathrm{Tr}_{p^{m}/p}(\kappa)+b\hskip 1.9919pt\mathrm{Tr}_{p^{m}/p}(\chi)$, for all $a,b\in\mathbb{F}_{p}$ and $\kappa,\chi\in\mathbb{F}_{p^{m}}$. We note that for an element $b\in\mathbb{F}_{p}$, $\mathrm{Tr}_{p^{m}/p}(b)=b$.
The group that generates the operator basis for $\mathbb{C}^{p^{m}\times p^{m}}$ defined in terms of the field based representation of basis states is [31]
$\mathcal{G}_{p^{m}}^{(g)}\!=\!\begin{cases}\left\{\omega^{l}\mathrm{X}^{(p^{m})}(\kappa)\mathrm{Z}^{(p^{m})}(\chi)\bigg{|}\kappa,\!\chi\in\mathbb{F}_{p^{m}}\text{ and }l\!\in\!\mathbb{Z}_{p}\right\},&\text{when characteristic $p$ is odd},\\ \left\{\mathrm{i}^{g}\omega^{l}\mathrm{X}^{(p^{m})}(\kappa)\mathrm{Z}^{(p^{m})}(\chi)\bigg{|}\kappa,\!\chi\in\mathbb{F}_{p^{m}}\text{ and }g,l\!\in\!\mathbb{Z}_{p}\right\},&\text{when characteristic $p$ is even},\\ \end{cases}$
where $\omega=\mathrm{e}^{\frac{\mathrm{i}2\pi}{p}}$, $\mathrm{i}=\sqrt{-1}$,
$\mathrm{X}^{(p^{m})}(\kappa)\ket{\theta}_{p^{m}}$ :=|$:=\ket{\kappa+\theta}_{p^{m}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall\theta\in\mathbb{F}_{p^{m}},$ (4)
$\mathrm{Z}^{(p^{m})}(\chi)\ket{\theta}_{p^{m}}$ $:=\omega^{\mathrm{Tr}_{p^{m}/p}(\chi\theta)}\ket{\theta}_{p^{m}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall\theta\in\mathbb{F}_{p^{m}}.$ (5)
We note that the factor $\mathrm{i}^{g}$ is included in the basis $\mathcal{G}_{p^{m}}$ when the characteristic $p$ is even as $\mathrm{iI}$ belongs to $\mathcal{P}$ and $\mathcal{P}=\mathcal{G}_{p^{m}}$ for $p=2$ and $m=1$.
The operator basis for $\mathbb{C}^{p^{m}\times p^{m}}$ is
$\mathcal{G}_{p^{m}}\!=\!\left\{\mathrm{X}^{(p^{m})}(\kappa)\mathrm{Z}^{(p^{m})}(\chi)\bigg{|}\kappa,\!\chi\in\mathbb{F}_{p^{m}}\right\},$ (6)
From equations (4) and (5), $\mathrm{X}^{(p^{m})}(\kappa)$ and $\mathrm{Z}^{(p^{m})}(\chi)$ are given by
$\mathrm{X}^{(p^{m})}(\kappa)$ $=\underset{\theta\in\mathbb{F}_{p^{m}}}{\sum}\ket{\kappa+\theta}\bra{\theta},$ (7)
$\mathrm{Z}^{(p^{m})}(\chi)$ $=\underset{\theta\in\mathbb{F}_{p^{m}}}{\sum}\omega^{\mathrm{Tr}_{p^{m}/p}(\chi\theta)}\ket{\theta}\bra{\theta}.$ | . (8)
The basis operator of the form $\mathrm{X}^{(p^{m})}(\kappa)\mathrm{Z}^{(p^{m})}(\chi)$ is uniquely represented by a vector of length $2$ defined over field $\mathbb{F}_{p^{m}}$, namely $[\kappa|\chi]_{p^{m}}$
$\mathrm{X}^{(p^{m})}(\kappa)\mathrm{Z}^{(p^{m})}(\chi)\equiv[\kappa|\chi]_{p^{m}}.$
The above defined operators will be used in Section 5.3 during the encoding and error correction procedures of our proposed quantum codes. In order to construct quantum error-correcting codes, we first derive a necessary and sufficient condition for $(\gamma,\Delta)$-cyclic codes to be dual containing. Note that a quantum code $[[n,k,d]]_{q}$ is said to be better than $[[n^{\prime},k^{\prime},d^{\prime}]]_{q}$ if any one of the following or both hold:
@@ 5.1 Quantum states and operators over qudits
Next, we define a Gray map and study $\mathbb{F}_{q}$-images of $(\gamma,\Delta)$-cyclic codes. Let $GL_{s+1}(\mathbb{F}_{q})$ be the set of all $(s+1)\times(s+1)$ invertible matrices over $\mathbb{F}_{q}$. Now, $\varphi:\mathscr{R}_{q,s}\longrightarrow\mathbb{F}_{q}^{s+1}$ define by
@@
Proof.
Let $a=(a_{0},a_{1},\dots,a_{n-1}),\leavevmode\nobreak\ b=(b_{0},b_{1},\dots,b_{n-1})\in\mathscr{R}_{q,s}^{n}$, where $a_{j}=\sum_{i=0}^{s}\zeta_{i}a_{j}^{i}$, $b_{j}=\sum_{i=0}^{s}\zeta_{i}b_{j}^{i}$ for $j=0,1,\dots,n-1$ and $a_{j}^{i},\leavevmode\nobreak\ b_{j}^{i}\in\mathbb{F}_{q}$. Then
@@
Moreover, $d_{G}(a,b)=\omega_{G}(a-b)=\omega_{H}(\varphi(a-b))=\omega_{H}(\varphi(a)-\varphi(b))=d_{H}(\varphi(a),\varphi(b))$. Hence, $\varphi$ is a distance preserving map. ∎
Theorem 5.1.
@@
Proof.
Follows directly from Proposition 5.1 and the definition of the Gray map. ∎
@@
Proof.
Let $c=(c_{0},c_{1},\dots,c_{n-1})\in\mathcal{C}$ and $d=(d_{0},d_{1},\dots,d_{n-1})\in\mathcal{C}^{\perp}$ where $a_{j}=\sum_{i=0}^{s}\zeta_{i}c_{j}^{i}$, $b_{j}=\sum_{i=0}^{s}\zeta_{i}d_{j}^{i}$ for $j=0,1,\dots,n-1$ and $a_{j}^{i},\leavevmode\nobreak\ b_{j}^{i}\in\mathbb{F}_{q}$. Now, $c\cdot d=\sum_{j=0}^{n-1}c_{j}d_{j}=0$ gives $\sum_{j=0}^{n-1}(c^{0}_{j}d^{0}_{j}+c^{1}_{j}d^{1}_{j}+\dots+c^{s}_{j}d^{s}_{j})=0$. Again,
@@
Since $c\in\mathcal{C}$ and $d\in\mathcal{C}^{\perp}$ are arbitrary, $\varphi(\mathcal{C}^{\perp})\subseteq(\varphi(\mathcal{C}))^{\perp}$. On the other hand, as $\varphi$ is a bijective linear map, $\mid\varphi(\mathcal{C}^{\perp})\mid=\mid(\varphi(\mathcal{C}))^{\perp}\mid$. Therefore, $\varphi(\mathcal{C}^{\perp})=(\varphi(\mathcal{C}))^{\perp}$. ∎
5.2 CSS Code Framework
Calderbank, Shor, and Steane [12] [44] proposed a framework to construct quantum error correction codes over qubits from two classical binary codes $C_{1}$ and $C_{2}$ that satisfy $C_{1}^{\perp}\subset C_{2}$. This class of codes are called the Calderbank-Shor-Steane (CSS) codes. The condition $C_{1}^{\perp}\subset C_{2}$ is called the dual-containing condition of CSS codes. By considering the two codes $C_{1}$ and $C_{2}$ to be the same code, i.e., $C_{1}=C_{2}$, we can construct quantum codes from dual-containing classical codes as $C_{1}^{\perp}\subset C_{2}=C_{1}$.
The CSS codes form a class of stabilizer codes. Let $H_{1}$ and $H_{2}$ be the parity check matrices of the classical codes $C_{1}[n,k_{1},d_{1}]$ and $C_{2}[n,k_{2},d_{2}]$, respectively. As $C_{1}^{\perp}\subset C_{2}$, the elements of $C_{1}^{\perp}$ are codewords of $C_{2}$; hence, $H_{2}H_{1}^{\mathrm{T}}=0$.
The CSS code is defined in the following two equivalent ways:
  • 1)
    The coset-based definition: As $C_{1}^{\perp}\subset C_{2}$, cosets of $C_{1}^{\perp}$ are formed in $C_{2}$. The basis codewords of the CSS code $\mathcal{Q}_{\mathrm{CSS}}$ are the normalized superposition of all the elements in a particular coset of $C_{1}^{\perp}$ in $C_{2}$. As $C_{1}^{\perp}$ has $2^{(n-k_{1})}$ elements and $C_{2}$ has $2^{k_{2}}$ elements, we obtain $C_{2}$ to contain $(2^{k_{2}})/(2^{(n-k_{1})})=2^{(k_{1}+k_{2}-n)}$ cosets of $C_{1}^{\perp}$. As each coset corresponds to a basis codeword, $\mathcal{Q}_{\mathrm{CSS}}$ has a dimension of $2^{(k_{1}+k_{2}-n)}$.
    Let $\omega_{0}$, $\omega_{1}$, $\dots$, $\omega_{g-1}$ be the cosets of $C_{1}^{\perp}$ in $C_{2}$, where $g=2^{(k_{1}+k_{2}-n)}$. Let $w_{0}$, $\dots$, $w_{g-1}$ be the coset representatives of the cosets $\omega_{0}$, $\dots$, $\omega_{g-1}$. The basis codeword of the CSS code corresponding to the coset $\omega_{i}$ ($i\in\{0,1,\dots,g-1\}$) is
    |$\ket{\psi_{i}}=\frac{1}{2^{((n-k_{1})/2)}}\sum_{l\in\mathcal{C}_{1}^{\perp}}\ket{l+w_{i}}.$ (9)
    The basis states in the superposition help to detect/correct the bit flip errors while the superposition helps to detect/correct the phase flip errors.
  • 2)
    The check matrix based definition: The check matrix of the CSS code [36] is
    $\mathcal{H}_{\mathrm{CSS}}=\left[\begin{array}[]{c|c}H_{1}&\mathbf{0}\\ \mathbf{0}&H_{2}\end{array}\right].$ (12)
The quantum codes obtained from both these definitions are the same for qubits.
Let $\rho_{1}=(n-k_{1})$ and $\rho_{2}=(n-k_{2})$. From equation (12), the first $\rho_{1}$ stabilizer generators that correspond to $[H_{1}|0]$ operate only the bit flip operator on a few qubits. They do not operate phase flip operators. As the bit flip and phase flip operators do not commute with each other, these stabilizers are used to detect and correct the phase flip errors. Similarly, the stabilizers that correspond to $[0|H_{2}]$ detect and correct the bit flip errors.
As the stabilizer code [18] correct bit flip errors and phase flip errors based on the stabilizers in $[0|H_{2}]$ and $[H_{1}|0]$, their bit flip and phase flip error correction capabilities are based on the error correction capabilities of $H_{2}$ and $H_{1}$, respectively. The minimum distance of the code is obtained to be $d^{\prime}\geq\mathrm{min}(d_{1},d_{2})$ [36].
Suppose that the parity check matrices $H_{1}$ and $H_{2}$ are full rank matrices. The check matrix in equation (12) is a $((\rho_{1}+\rho_{2})\times 2n)$ matrix. As $H_{1}$ and $H_{2}$ are full rank matrices, the CSS code has $(\rho_{1}+\rho_{2})$ minimal stabilizer generators. Thus, the size of the CSS code is $2^{(n-(\rho_{1}+\rho_{2}))}=2^{(k_{1}+k_{2}-n)}$. Hence, the CSS code is an $[[n,k_{1}+k_{2}-n,d^{\prime}\geq\mathrm{min}(d_{1},d_{2})]]$ stabilizer code.
Next, we discuss the CSS code over qudits that is obtained from the classical codes $D_{1}$ and $D_{2}$ by using two different approaches for obtaining the basis codewords.
  1. 1.
    Coset-based construction of the CSS code: As $D_{1}^{\perp}$ is a subset of $D_{2}$, there exist cosets of $D_{1}^{\perp}$ in $D_{2}$. The size of $D_{1}^{\perp}$ and $D_{2}$ are $p^{m(n-k_{1})}$ and $p^{mk_{2}}$, respectively; hence, the number of cosets of $D_{1}^{\perp}$ in $D_{2}$ is $s^{\prime}=(p^{mk_{2}}/p^{m(n-k_{1})})=p^{m(k_{2}-n+k_{1})}=p^{m(k_{1}+k_{2}-n)}$. Thus, the dimension of the quantum code obtained is $p^{m(k_{1}+k_{2}-n)}$, similar to the CSS code over qubits whose dimension is $2^{(k_{1}+k_{2}-n)}$.
    Let $\tau_{0}$, $\tau_{1}$, $\dots$, $\tau_{(s^{\prime}-1)}$ be the $s^{\prime}$ cosets of $D_{1}^{\perp}$ in $D_{2}$. Let ${t_{0}}$, ${t_{1}}$, $\dots$, ${t_{(s^{\prime}-1)}}$ be the coset representatives of $\tau_{0}$, $\tau_{1}$, $\dots$, $\tau_{(s^{\prime}-1)}$, respectively. The basis codeword |$\ket{\psi_{i}^{(p^{m})}}$ ($i\in\{0,1,\dots,s^{\prime}-1\}$) of the CSS code over qudits obtained from the coset $\tau_{i}$ is
    |$\ket{\psi_{i}^{(p^{m})}}=\frac{1}{p^{m((n-k_{1})/2)}}\sum_{l\in\mathcal{D}_{1}^{\perp}}\ket{l+t_{i}}.$ (13)
  2. 2.
    Parity check matrix of the CSS code ([35]) : The check matrix for the CSS code obtained from $D_{1}$ and $D_{2}$ that satisfy $D_{1}^{\perp}\subset D_{2}$, whose basis codewords are provided in Equation 13, is given by,
    $\mathcal{H}_{\mathrm{CSS}}^{(p^{m})}=\left[\begin{array}[]{c|c}\begin{matrix}H_{d_{1}}\\ \alpha H_{d_{1}}\\ \vdots\\ \alpha^{m-1}H_{d_{1}}\end{matrix}&\mathbf{0}\\ \mathbf{0}&\begin{matrix}H_{d_{2}}\\ \alpha H_{d_{2}}\\ \vdots\\ \alpha^{m-1}H_{d_{2}}\end{matrix}\end{array}\right],$ (16)
    where $\alpha$ is the primitive element of $\mathbb{F}_{p^{m}}$.
Now, keeping the above discussion in mind, we derive a necessary and sufficient condition for dual-containment. Currently, CSS construction (Lemma 5.2) is one of the widely used techniques to obtain quantum codes from classical linear codes, in which dual containing linear codes play an instrumental role.
Lemma 5.2 ([19], Theorem 3).